OpenStudy (anonymous):

Find the thirty-second term of the following sequence. 9, 15, 21, ...

OpenStudy (amoodarya):

9, 15, 21, ... 9 ,9+6 ,9+2(6) ,... so a1=9 a2=15 d=15-9=21-15=6 so a(n)a1+(n-1)d a(n)=9+6(n-1) so now put n=32

OpenStudy (anonymous):

i dont get it man